OpenStudy (anonymous):

f(x)=-7+2x,what is f(8)

OpenStudy (dmndlife24):

Plug in 8 for x... f(8) = -7 + 2(8) f(8) = ?

OpenStudy (anonymous):

f(8) = -7+2(8) f(8) = -5(8) f(8) = -10 Answer: f(8) = -10